OpenStudy (anonymous):

Evan read 10 pages of a book. He plans to read 8 pages a day until he finishes the book. The expression below represents the total pages read,where d is the number of days.How many pages will Evan have read after 9 days? 8d + 10

OpenStudy (anonymous):

Put the number 9 in for d and then use order of operations to get the value.

OpenStudy (anonymous):

8d + 10 = p p is the number of pages read here we insert 9 for d. 8(9) + 10 = p 62 + 10 = p 72 = p He read 72 pages

OpenStudy (anonymous):

thanks